OpenStudy (anonymous):

Which of the following exponential functions goes through the points (1, 6) and (2, 12)?

OpenStudy (solomonzelman):

Ok, an exponential function is generally written as: \(\large\color{black}{ \displaystyle y=a(b)^x }\) You have two (x,y) coordinates given, and them you shall use to solve for a and b.

OpenStudy (solomonzelman):

With these two points, you will set the following system of eqations up to solve for a and b. \(\large\color{black}{ \displaystyle 6=a(b)^1 }\) \(\large\color{black}{ \displaystyle 12=a(b)^2 }\)

OpenStudy (solomonzelman):

Is this making some sense?